# Lax dynamics

### v1

@@ Abstract
pacs:
Introduction – The solution of a physical problem usually proceeds by identifying a solvable part and studying the effects of perturbations. For nonlinear systems where the solvable part is described by an integrable classical or quantum Hamiltonian, one can, with varying degrees of mathematical difficulty, separate the independent degrees of freedom (the quasiparticles) and analyze their interactions using, for example, perturbation theory. In many-body problems, a perturbation typically destroys integrability, leaving only a few residual conserved quantities, it is important to assess how and when thermalization, chaotic dynamics, and conventional hydrodynamic behavior occur [1]. Considering that a variety of physical systems as ultracold atoms, one-dimensional magnets, or optical beams are proximate to nonlinear integrable limits [2], such questions are of wide interests in many diverse contexts.
@@
$L\ket{\alpha}=\lambda_{\alpha}\ket{\alpha},\quad\alpha=1,\dots,N.$ (6)
where $\lambda_{\alpha}$ and |$\ket{\alpha}$ are the Lax eigenvalues and eigenvectors, whose components in the lattice basis |$\ket{j}$ are $\psi_{\alpha,j}=\braket{\alpha}{j}$.
Integrability - For Toda ($U=0$), $L(t)$ is isospectral i.e. $\lambda_{\alpha}$ do not change while eigenvectors are time-dependent and satisfy |$\dot{\ket{\alpha}}=B\ket{\alpha}.$ At zero temperature the spectrum is $\lambda_{\alpha}=2\cos(2\pi\alpha/N)$ and eigenstates with $|\lambda_{\alpha}|>1$ are associated to solitons [27]. The physically relevant conserved charges are given by $\mathrm{tr}\big{(}L^{n}\big{)}=\sum_{j=1}^{N}(L^{n})_{j,j}=\sum_{\alpha=1}^{N}\lambda_{\alpha}^{n}$. The first three are the standard ones, namely the sum of the stretches $r_{j}$, momenta $p_{j}$ and local energies $p_{j}^{2}+a_{j}^{2}+a_{j-1}^{2}$.
@@
Quasiparticles – The quasiparticle concept is insightful to understand the dynamics [10, 29, 30, 31]. For integrable systems quasiparticles move ballistically. Upon collisions they retain their velocity but undergo a spatial shift, the case termed interacting in Ref. [32]. This results in an effective velocity, which for Toda is solution of suitable integral equation [9]. To visualize this concept, one can define the quasiparticle position and velocity $x_{\alpha},v_{\alpha}$ as the center of mass of the Lax eigenvector [30]
@@
Conservative perturbations of the Toda chain – Upon multiplying both sides of Eq.(5) by $L^{n-1}$, and using cyclic and linearity properties of the trace
@@
Examples – In principle, for any Hamiltonian $H$ one can recast the equation of motion in the form (5). However, for a generic perturbation, $U$ may have a complicated dependence on the $L_{i,j}$: for instance, terms like $r_{j}^{p}$ in $\Phi$ would yield entries proportional to $\log^{p-1}a_{j}$ [26]. It is thus useful to examine some simpler cases. The first is the Morse chain $\Phi(x)=(e^{-x/2}-\varepsilon)^{2}$ [33] then (up to a constant)
@@
Lax dynamics – The main idea is now to look at the time evolution of the $\lambda_{\alpha}$ that, under the effect of the perturbation $U$, are no longer constant  [^2] To the best of my knowledge, this point of has been considered only very sporadically in the literature, see for example the case of the driven-dissipative Toda chain [46] and Toda with a point-like defect [47]. . For illustration, let us compare simulations of the Toda and Morse chains. In agreement with intuition, the space-time evolution of a Lax eigenvector $|\psi_{\alpha,n}(t)|^{2}$ [Fig. 1(a)] looks soliton-like in the integrable case, propagating ballistically with random space shifts, see Fig. 1(b) where the quasi-particle velocity $v_{\alpha}(t)$ is reported along with the corresponding (constant) $\lambda_{\alpha}$. Plotting $\lambda_{\alpha}$ versus the time-averaged velocity $\langle v_{\alpha}\rangle$ [Fig. 1(c)] confirms an approximate correspondence between the two quantities, indicating that one can be used as a proxy of the other. Also, they become almost exactly proportional for soliton modes located closer to the Lax spectrum band edges.
@@
For an analytical formulation of Lax dynamics, one proceeds by computing $d(L\ket{\alpha})/dt$ from Eq. (6) and using Eq. (5), yielding
$U\ket{\alpha}+(L-\lambda_{\alpha})\big{(}\dot{\ket{\alpha}}-B\ket{\alpha}\big{)}=\dot{\lambda}_{\alpha}\ket{\alpha}.$ (11)
Upon left-multipying by $\bra{\beta}$ |, using $\braket{\alpha}{\beta}=\delta_{\alpha,\beta}$, and letting $\braket{\alpha}{U}{\beta}\equiv U_{\alpha\beta}$ one obtains from the diagonal elements that $\dot{\lambda}_{\alpha}=U_{\alpha\alpha}$ and from the non-diagonal elements the evolution equation for the eigenvectors
|$\ket{\dot{\alpha}}=B\ket{\alpha}\,+\,\sum_{\beta\neq\alpha}\frac{U_{\alpha\beta}}{\lambda_{\alpha}-\lambda_{\beta}}\ket{\beta}$ (12)
@@
where $F_{\alpha\beta}\equiv\braket{\alpha}{\left(\dot{U}+[U,B]\right)}{\beta}$. Eqs. (Lax dynamics) are exact and hold whatever the form of $U$. Also, using the invariance of the trace with respect to change of basis, if momentum conservation holds for also in presence of perturbation $trU=0$, the the ”momentum”, $\sum_{\alpha}\pi_{\alpha}$ is conserved too.
If $F_{\alpha\beta}$ would vanish, Eqs. (Lax dynamics) would be identical to the so-called Pechukas-Yukawa (PY) equations ruling the evolution of the eigenenergies of a quantum Hamiltonian of the form $\mathcal{H}=\mathcal{H}_{0}+t\mathcal{H}_{1}$, as a function of the fictitious ”time” parameter $t$ [37, 38]. Also, Eq. (12) would formally correspond to those for quantum eigenstates [39]. The PY equations are closed and define a $22N+N(N-1)/$ 2-dimensional dynamical system ($U$ is symmetric) with generalized coordinates $(\lambda_{\alpha},\pi_{\alpha},U_{\alpha\beta})$. It has the remarkable features to be both Hamiltonian and fully integrable. Indeed, the change of variable $V_{\alpha\beta}={U_{\alpha\beta}}{(\lambda_{\alpha}-\lambda_{\beta})}$ ($\alpha\neq\beta$) transforms the PY equations in a generalized version of the famous Calogero-Moser model [39, 40].
Yet, the presence of the terms $F_{\alpha\beta}$ hinders an exact analysis as in the PY case: indeed, Eqs. (Lax dynamics) are not closed and, in general, are also explicitly time-dependent  [^4] Indeed also in the quantum problem, a different (nonlinear) dependence of the Hamiltonian $\mathcal{H}(t)$ on the parameter $t$ may yield additional non-autonomous terms, that can be handled analytically only is specific cases, see e.g. Ref.[48]. Despite those difficulties, some interesting physical consequences can be envisaged.
First, ignoring for the moment the terms $F_{\alpha\alpha}$, the first two of Eqs.(Lax dynamics) describe a one-dimensional Dyson-Coulomb gas coupled through the fluctuating ”charges” $U_{\alpha\beta}^{2}$ provided by the remaining $N(N-1)/2$ degrees of freedom that act as a ”heat bath” [42]. Indeed, for a finite number of levels, the DOS is approximatively given by the the invariant measure of the Coulomb gas under an external quadratic potential [38]. On the other hand, for the Toda chain, such measure coincides also with the thermal DOS $\rho_{th}$ [10]. Indeed, based on the above heuristic consideration, one may surmise that the Lax dynamics naturally provides a general thermalization pathway towards the thermal DOS for any perturbation in the class of the above examples. In other words, the fluctuations of the charges provide the chaos (noise) source needed to thermalize the Dyson gas. It may be argued, that the generic mechanism leading to thermalization is provided by level repulsion and quasi-elastic scattering seen in Fig.2.
@@
Conclusions – Lax dynamics is a novel and insightful approach to describe the effect of perturbations on a many-body integrable system at finite energy density. Its implementation is computationally straightforward, and it has the potential to be extended to other systems, such as perturbations of integrable discretizations of the nonlinear Schrödinger equation [43], among others. It offers a physically appealing interpretation in terms of quasi-particle collisions as avoided eigenvalue crossings. Furthermore, in the case of weak perturbations, Lax dynamics is expected to evolve on a slower timescale compared to the natural timescale of the Flashka variables. Therefore, it could be an effective approach for studying the slow evolution of Toda actions in the context of thermalization problems [16, 19]. Finally, the analogy with the PY gas is highly suggestive and warrants a more detailed investigation. Equations (Lax dynamics) may allow for an effective, reduced dynamical description of the relevant quantities. The ’universal’ evolution described by PY dynamics is one of the arguments used to justify the universal spectral statistics of quantum chaos. Could a similar consideration also apply in the present context? This could be one of the many possible research routes originating from the present work.
Acknowledgements – Support is acknowledged from the PRIN 2022 project Breakdown of ergodicity in classical and quantum many-body systems (BECQuMB) Grant No. 20222BHC9Z CUP G53C24000680006 funded by the European Union— NextGenerationEU, M4 C2 1.1. I thank L. Cugliandolo, S.Iubini, A. Politi and S. Ruffo for useful discussions.

### v2

# Istituto Nazionale di Fisica Nucleare, Sezione di Firenze, via G. Sansone 1, I-50019 Sesto Fiorentino, Italy

@@ Abstract
1 Introduction
The solution of a physical problem usually proceeds by identifying a solvable part and studying the effects of perturbations. For nonlinear systems where the solvable part is described by an integrable classical or quantum Hamiltonian, one can, with varying degrees of mathematical difficulty, separate the independent degrees of freedom (the quasiparticles) and analyze their interactions using, for example, perturbation theory. In many-body problems, a perturbation typically destroys integrability, leaving only a few residual conserved quantities, it is important to assess how and when thermalization, chaotic dynamics, and conventional hydrodynamic behavior occur [1]. Considering that a variety of physical systems as ultracold atoms, one-dimensional magnets, or optical beams are proximate to nonlinear integrable limits [2], such questions are of wide interests in many diverse contexts.
@@ 1 Introduction
[…]
2 Integrability
[…]
@@ 2 Integrability
3 Quasiparticles
The quasiparticle concept is insightful to understand the dynamics [10, 29, 30, 31]. For integrable systems quasiparticles move ballistically. Upon collisions they retain their velocity but undergo a spatial shift, the case termed interacting in Ref. [32]. This results in an effective velocity, which for Toda is solution of suitable integral equation [9]. The idea that these quasiparticles, acquire an effective, constant velocities has been also rigorously established [33]. There is also an intriguing connection between such concepts and that of Anderson localization, rooted in the basically random nature of Lax matrices [34].
To visualize this concept, one can define the quasiparticle position and velocity $x_{\alpha},v_{\alpha}$ as the center of mass of the Lax eigenvector [30]
@@ 3 Quasiparticles
4 Conservative perturbations of the Toda chain
Upon multiplying both sides of Eq.(5) by $L^{n-1}$, and using cyclic and linearity properties of the trace
@@ 4 Conservative perturbations of the Toda chain
5 Examples
In principle, for any Hamiltonian $H$ one can recast the equation of motion in the form (5). However, for a generic perturbation, $U$ may have a complicated dependence on the $L_{i,j}$: for instance, terms like $r_{j}^{p}$ in $\Phi$ would yield entries proportional to $\log^{p-1}a_{j}$ [26]. It is thus useful to examine some simpler cases. The first is the Morse chain $\Phi(x)=(e^{-x/2}-\varepsilon)^{2}$ [35] then (up to a constant)
@@ 5 Examples
6 Lax dynamics
The main idea is now to look at the time evolution of the $\lambda_{\alpha}$ that, under the effect of the perturbation $U$, are no longer constant  [^2] To the best of my knowledge, this point of has been considered only very sporadically in the literature, see for example the case of the driven-dissipative Toda chain [38] and Toda with a point-like defect [39]. . For illustration, let us compare simulations of the Toda and Morse chains. In agreement with intuition, the space-time evolution of a Lax eigenvector $|\psi_{\alpha,n}(t)|^{2}$ [Fig. 1(a)] looks soliton-like in the integrable case, propagating ballistically with random space shifts, see Fig. 1(b) where the quasi-particle velocity $v_{\alpha}(t)$ is reported along with the corresponding (constant) $\lambda_{\alpha}$. Plotting $\lambda_{\alpha}$ versus the time-averaged velocity $\langle v_{\alpha}\rangle$ [Fig. 1(c)] confirms an approximate correspondence between the two quantities, indicating that one can be used as a proxy of the other. Also, they become almost exactly proportional for soliton modes located closer to the Lax spectrum band edges.
@@ 6 Lax dynamics
[…]
@@
where $F_{\alpha\beta}\equiv\braket{\alpha|\left(\dot{U}+[U,B]\right)|\beta}$. Eqs. (6) are exact and hold whatever the form of $U$. Also, using the invariance of the trace with respect to change of basis, if momentum conservation holds for also in presence of perturbation $trU=0$, the the ”momentum”, $\sum_{\alpha}\pi_{\alpha}$ is conserved too.
If $F_{\alpha\beta}$ would vanish, Eqs. (6) would be identical to the so-called Pechukas-Yukawa (PY) equations ruling the evolution of the eigenenergies of a quantum Hamiltonian of the form $\mathcal{H}=\mathcal{H}_{0}+t\mathcal{H}_{1}$, as a function of the fictitious ”time” parameter $t$ [40, 41]. Also, Eq. (12) would formally correspond to those for quantum eigenstates [42]. The PY equations are closed and define a 2N+N(N1)/22N+N(N-1)/2-dimensional dynamical system ($U$ is symmetric) with generalized coordinates $(\lambda_{\alpha},\pi_{\alpha},U_{\alpha\beta})$. It has the remarkable features to be both Hamiltonian and fully integrable. Indeed, the change of variable $V_{\alpha\beta}={U_{\alpha\beta}}{(\lambda_{\alpha}-\lambda_{\beta})}$ ($\alpha\neq\beta$) transforms the PY equations in a generalized version of the famous Calogero-Moser model [42, 43].
Yet, the presence of the terms $F_{\alpha\beta}$ hinders an exact analysis as in the PY case: indeed, Eqs. (6) are not closed and, in general, are also explicitly time-dependent  [^4] Indeed also in the quantum problem, a different (nonlinear) dependence of the Hamiltonian $\mathcal{H}(t)$ on the parameter $t$ may yield additional non-autonomous terms, that can be handled analytically only is specific cases, see e.g. Ref.[44]. Despite those difficulties, some interesting physical consequences can be envisaged.
First, ignoring for the moment the terms $F_{\alpha\alpha}$, the first two of Eqs.(6) describe a one-dimensional Dyson-Coulomb gas coupled through the fluctuating ”charges” $U_{\alpha\beta}^{2}$ provided by the remaining N(N1)/2N(N-1)/2 degrees of freedom that act as a ”heat bath” [45]. Indeed, for a finite number of levels, the DOS is approximatively given by the the invariant measure of the Coulomb gas under an external quadratic potential [41]. On the other hand, for the Toda chain, such measure coincides also with the thermal DOS $\rho_{th}$ [10]. Indeed, based on the above heuristic consideration, one may surmise that the Lax dynamics naturally provides a general thermalization pathway towards the thermal DOS for any perturbation in the class of the above examples. In other words, the fluctuations of the charges provide the chaos (noise) source needed to thermalize the Dyson gas. It may be argued, that the generic mechanism leading to thermalization is provided by level repulsion and quasi-elastic scattering seen in Fig.2.
@@
7 Conclusions
Lax dynamics is a novel and insightful approach to describe the effect of perturbations on a many-body integrable system at finite energy density. Its implementation is computationally straightforward, and it has the potential to be extended to other systems, such as perturbations of integrable discretizations of the nonlinear Schrödinger equation [46], among others. It offers a physically appealing interpretation in terms of quasi-particle collisions as avoided eigenvalue crossings. Furthermore, in the case of weak perturbations, Lax dynamics is expected to evolve on a slower timescale compared to the natural timescale of the Flashka variables. Therefore, it could be an effective approach for studying the slow evolution of Toda actions in the context of thermalization problems [16, 19]. Finally, the analogy with the PY gas is highly suggestive and warrants a more detailed investigation. Equations (6) may allow for an effective, reduced dynamical description of the relevant quantities. The ’universal’ evolution described by PY dynamics is one of the arguments used to justify the universal spectral statistics of quantum chaos. Could a similar consideration also apply in the present context? This could be one of the many possible research routes originating from the present work.
Acknowledgements.
Support is acknowledged from the PRIN 2022 project Breakdown of ergodicity in classical and quantum many-body systems (BECQuMB) Grant No. 20222BHC9Z CUP G53C24000680006 funded by the European Union— NextGenerationEU, M4 C2 1.1. I thank L. Cugliandolo, S.Iubini, A. Politi and S. Ruffo for useful discussions. Data availability statement: The data that support the findings of this study are available upon reasonable request from the author.